OpenStudy (anonymous):

y=3X+1 What is the domain and the range of this question?

OpenStudy (anonymous):

domain is xis element of all real numbers and range is similar there are no points of discontinuity

OpenStudy (anonymous):

\[\left( \infty,-\infty \right)\] in interval notation. OR all real numbers